OpenStudy (anonymous):

x^2-4x+2=0 using the quadratic fomula solve this

OpenStudy (anonymous):

x=-b+-sq rt b^2-4ac/2a ax^2+bx+c=0

OpenStudy (anonymous):

so a=1, b=4 and c=2 plug into the first equation and solve

OpenStudy (anonymous):

\[x=b \pm \sqrt{b ^{2-4ac}}/2a\]

OpenStudy (anonymous):

plug the values into the previous equation and solve

OpenStudy (anonymous):

just so the equation is a little clearer: \[x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\frac{numerator}{denominator}

OpenStudy (anonymous):

what did you get as an anwer

OpenStudy (anonymous):

answer sorry

OpenStudy (anonymous):

use joe maths equation it is clearer and plug in so a=1, b=4 and c=2

OpenStudy (anonymous):

-3.414213562, -0.5857864376 are your answers